Given the linear equation y = -2x – 4, find the y-coordinates of the points (0, ), (-1, ), and (2, ). Please show all of your work. Plot these points and graph the linear equation. Submit your graph through the dropbox.
Answer by mananth(16949)

y=-2x-4
(0, )
y=-2*0-4
y=-4
(0,-4)

(-1, )
y=-1*-2-4
y=-2
(-1,-2)
, and (2, ).
y=2*-2-4
y=-8
(2,-8)